How would you solve this equation?
x^2-49=48x

Answer by algebrapro18(206)

first get the expression on one side
x^2-48x-49=0

now factor
(x-49)(x+1)=0

now set each factor equal to 0
x-49 =0 x+1=0

solve
x = 49 and x = -1